On a certain day in Glacier Park, 728 eagles were counted. Five years later, 594 were counted. What was the percentage of decrease in the number of eagles counted?
Please help, thank you!

Answer by checkley77(12844)

(728-594)/728=134/728=.184 or 18.4% decrease.